OpenStudy (anonymous):

simplify the expression 5 square root 24 * 3 square root 10

OpenStudy (johnweldon1993):

\[\large 5\sqrt{24} \times 3\sqrt{10}\] Hint...when we multiply radicals...we just multiply under the radical Example \[\large \sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\] so here, multiply the 5 and the 3...and then also multiply the sqrt(24) and sqrt(10)

OpenStudy (johnweldon1993):

Well it would actually be *remember the 24 and the 10 stay under the radical...so we have \[\large 15\sqrt{240}\] Can that be simplified at all? for exampe if I wrote this as \[\large 15 \sqrt{16 \times 15}\] anything can be simplified?

OpenStudy (johnweldon1993):

*Hint \[\large \sqrt{16} = 4\]

OpenStudy (anonymous):

15 cant be simplified so is it 4* 15 15 radical 60 ?